A car leaves town traveling at 50 miles per hour. Two hours later, a second car, traveling 70 miles per hour, leaves along the same route as the first car. How long would it take for the second car to overtake the first car?
I have tried to set up my rate distance time table but it just doesn't seem to be working for me. thank you very much for the help.

Answer by mananth(16946)

Car A --------------- 50mph
Car B -----------70 mph
The distance traveled by both is same when they meet
but Car A takes 2 hours more
Time car B = Time Car A +2
t=d/r
x/70 =x/50 -2
x/50 -x/70 =2
70x-50x= 2*50*70
20x=2*50*70
x= 2*50*70/20
x=50*7
x=350 miles
speed of Car B = 70 mph
distance = 350 miles
t= 350/70
t=5 hours
car B takes 5 hours